# MOKA: Open-Vocabulary Robotic Manipulation through Mark-Based Visual Prompting

### v1

@@ IV-A Motion with Point-based Affordances
To leverage VLMs for solving open-vocabulary manipulation tasks, there needs to be an interface that connects the inputs and outputs of the VLM and the motions performed by the robot. To achieve this goal, we design an affordance representation defined on 2D images. Produced as the end result by the VLM, the affordance representation specifies the desired motion.
@@
By extending the definitions in Manuelli et al. [32] and Qin et al. [38], we design a point-based affordance representation for a wide range of manipulation tasks. Instead of separately devising motion primitives for different pre-defined skills, we use a unified set of keypoints and waypoints to specify the motion. These points are predicted by VLMs on 2D images and converted to poses in the $\mathbf{SE(3)}$ space. Then a smooth motion trajectory is generated based on these poses. To perform the task, the robot gripper interacts with the environment by following the generated motion trajectory.
@@
We specify the robot’s motion in an object-centric manner as shown in Fig. 2. Following the discussion in Sec. III, we would like this representation to be applicable to different types of interactions with objects in the environment. Therefore, we consider two types of objects, $o_{\text{in-hand}}$ (e.g., the broom) and $o_{\text{unattached}}$ (e.g., the trash), and specify the motion with a grasping phase and a manipulation phase. In the grasping phase, the robot reaches and grasps an object $o_{\text{in-hand}}$ from the environment. Then in the manipulation phase, the robot performs a motion and makes contact with another object $o_{\text{unattached}}$, either directly or using $o_{\text{in-hand}}$ as a tool. In some scenarios, only one of these two types of objects is interacted with by the robot, either $o_{\text{in-hand}}$ (e.g., unplugging a cable, opening a drawer) or $o_{\text{unattached}}$ (e.g., pressing a button), and one of the two phases can be skipped accordingly.
@@
We now describe the definition of the keypoints and waypoints as well as how they are used to specify the motions in both phases. These points are illustrated in Fig. 2 and more examples can be found in Sec. V-C. Following the practice of Manuelli et al. [32] and Qin et al. [38], we use the grasping keypoint $x_{\text{grasp}}$ to specify the position on $o_{\text{in-hand}}$ where the robot gripper should hold the object. If $o_{\text{in-hand}}$ is not involved in a task, the grasping phase will be skipped. For the manipulation phase, the robot’s gripper follows a motion trajectory specified by an additional set of points. The function keypoint $x_{\text{function}}$ specifies the part of $o_{\text{in-hand}}$ that will make contact with $o_{\text{unattached}}$ in the manipulation phase. If $o_{\text{in-hand}}$ is not specified, $x_{\text{function}}$ will be on the robot gripper and the contact will directly be made between the robot and $o_{\text{unattached}}$. Correspondingly, the target keypoint $x_{\text{target}}$ is the part of $o_{\text{unattached}}$ that will be contacted by $x_{\text{function}}$ during the manipulation phase. We also introduce the pre-contact waypoints $x_{\text{pre-contact}}$ and the post-contact waypoints $x_{\text{post-contact}}$ defined in free space, which dictates the manipulation motion along with the keypoints defined on the objects.
@@
During the manipulation phase, the robot moves the gripper such that $x_{\text{function}}$ follows the path sequentially connecting the $x_{\text{pre-contact}}$, $x_{\text{target}}$, and $x_{\text{post-contact}}$. Besides following the path, we also require the robot gripper to follow the specified grasping orientation $R_{\text{grasp}}$ and manipulation orientation $R_{\text{manipulate}}$ during the two phases respectively. To better illustrate the design of our point-based motion, we provide examples of the predicted point specifications and the resultant motions from our experiments in Appendix and Sec. V-C. In MOKA, this set of keypoints and additional attributes (described in Sec. IV-C) are summarized in a dictionary as the affordance representation (see Fig. 2).
@@ IV-B Affordance Reasoning with Vision-Language Models
To predict the defined affordance representations, we employ the VLM $\mathcal{M}(\cdot)$ (defined in Sec. III), which is pre-trained on Internet-scale data for solving general visual question answering (VQA) problems. Using a hierarchical prompting framework as shown in Fig. 2, MOKA converts this affordance reasoning problem into a series of VQA problems that are solvable by the pre-trained VLM.
@@
The hierarchical prompting framework takes as input the free-form language description $l$ of the task and an RGB image observation of the environment $s_{t}$. MOKA examines the initial observation $s_{0}$ and decomposes the task $l$ into a sequence of subtasks using the VLM. For each of the subtasks, the VLM is asked to provide the summary of the subtask instruction, the description of the corresponding $o_{\text{in-hand}}$, $o_{\text{unattached}}$, as well as the description of the motion (e.g., ”from left to right”). On the low level, given the response from the high-level reasoning and the visual observation $s_{t(k)}$ at the beginning of $k$-th subtask (at the time step $t(k)$), the VLM is queried again with a different prompt to produce the affordance representation defined in Sec. IV-A. In the remainder of this section, we will describe the input formats, output formats, and prompt designs that we use to instantiate this method. Further details, including the complete prompts, can be found in the Appendix B.
@@
High-level reasoning. Given the initial observation $s_{0}$ and the language description $l$, we first query the VLM $\mathcal{M}$ with the language prompt $p_{\text{task}}$ to produce the response $y_{\text{high}}$:
$y_{\text{high}}=\mathcal{M}([p_{\text{high}},l,s_{0}]).$ (1)
@@
The representation $y_{\text{high}}$ is a string that contains structured information for the $K$ subtasks that the VLM infers are needed to solve the task. We design the prompt so as to require the VLM to produce $y_{\text{task}}$ as a list of dictionaries. As shown in Fig. 2, each dictionary contains the language description of a subtask (e.g., “Wipe the snack package to the right side of the table using the broom.”), as well as detailed information to facilitate motion generation, including the description of $o_{\text{in-hand}}$ (e.g., “broom”), the object name of $o_{\text{unattached}}$ (e.g., “snack package”), and the description of the motion (e.g., “from left to right”). This high-level plan will be used as an intermediate result for producing the detailed affordance representation through the low-level reasoning with the VLM.
@@
Low-level reasoning. Next, we prompt the VLM once again to produce the affordance representation defined in Sec. IV-A as $y_{\text{low}}^{k}$, conditioning on the high-level representation $y_{\text{high}}$ and the visual observation $s_{t(k)}$ at the beginning of the $k$-th subtask. Instead of directly predicting 3D coordinates on 2D images, which is challenging and even ill-defined, we query the VLM to output 2D coordinates on the images and deproject them back to the 3D space. The three keypoints $x_{\text{grasp}},x_{\text{function}}$ and $x_{\text{target}}$ are defined on the object surface, and thus we can compute the 3D coordinates using the corresponding depth value of the 2D location based on the RGB image and camera parameters. For the waypoints in free space, we query the VLM to predict the desired height in text. To produce such an affordance representation $y_{\text{motion}}^{k}$, we query the VLM again by
$y_{\text{low}}^{k}=\mathcal{M}([p_{\text{low}},y_{\text{task}}^{k},f(s_{t(k)})]),$ (2)
where $y_{\text{high}}^{k}$ is the substring corresponding to the $k$-th subtask extracted from $y_{\text{high}}$, and $f(\cdot)$ is a function that process the raw visual observation $s_{t(k)}$. We will explain the motivation and detailed implementation of $f(\cdot)$ in the next section and Appendix B. Through our ablation study in Appendix C, the hierarchical prompting strategy is essential for VLM to successfully perform the affordance reasoning for solving the tasks.
@@ IV-C Mark-Based Visual Prompting
Inspired by Yang et al. [52], MOKA uses a set of marks as visual prompts to enable VLM to apply its reasoning capability to predict the point-based affordance representation as shown in Fig. 2. Consisting of dots, grids, and text notations annotated on the image observation, these marks play an important role in the reasoning process. Proposed by open-vocabulary object detection and segmentation algorithms, these marks facilitate visual reasoning by encouraging the VLM to attend to the target objects and other task-relevant information in the image. We annotate marks as candidate parts and regions for the VLM to choose the points from, converting the original problem of directly generating coordinates into multiple-choice questions, which is usually more tractable for existing VLMs.
@@
To select keypoints, which are defined on the in-hand object $o_{\text{in-hand}}$ and the unattached object $o_{\text{unattached}}$ suggested by the high-level reasoning in Sec. IV-B, we propose and plot candidate keypoints on these objects. Given the names of $o_{\text{in-hand}}$ and $o_{\text{unattached}}$, we first segment these two objects using GroundedSAM [43], which combines GroundingDINO [27] and SAM [58] to extract segmentation masks of objects specified by a text prompt. After we obtain the segmentation masks of $o_{\text{in-hand}}$ and $o_{\text{unattached}}$, we perform farthest point sampling [37] on the object contour to obtain $K$ boundary points. Together with its geometric center, and overlay the $K+1$ candidate keypoints on each object. Each candidate keypoint is assigned an index, which is annotated next to it as a reference. To avoid confusion, we use different colors for candidate keypoints on $o_{\text{in-hand}}$ and $o_{\text{unattached}}$ and use the caption in the format of $P_{i}$ and $Q_{j}$ respectively, where $i$ and $j$ are integers. More implementation details can be found in Appendix B-B.
@@
Selecting waypoints in free space involves searching over a much larger region. Instead of directly sampling points in the entire workspace, we divide the observed RGB image into an $M\times n$ grid, where $m$ and $n$ are integers. Both $m$ and $n$ are set to $5$ for our evaluation tasks. The VLM is prompted to choose the tiles in which the pre-contact and post-contact keypoints are supposed to locate in and then the exact waypoints are sampled uniformly within the tile. For this purpose, we overlay the grid along with the name of each tile on the image. The tile names follow chess notation, which uses letters to specify the columns and integers for the rows.
IV-D Zero-shot Execution in Real World
@@
After we obtain the selected keypoints and waypoints from the VLM, we will convert the affordance to an executable motion on a real robot. To achieve this goal, we need to lift up all the points from the 2D image to 6D Cartesian space.
@@
For the keypoints defined on the object, we can directly take the corresponding point on the registered depth image and transform it into the robot’s frame. For the waypoints that are selected from the free space, since it’s not attached to any objects, its height needs to be specified in order to be deprojected to 3D space. For this purpose, we also ask the VLM to determine the height of waypoints. Here We only consider the cases where the waypoints are at the same height as the target point (e.g. pushing), or above the target point (e.g. placing) in most common tabletop manipulation scenarios. Additionally, we also prompt the VLM to return the orientation of $o_{\text{in-hand}}$ during the manipulation phase, which would usually affect the success of tool use tasks.
@@
Drawing inspirations from Manuelli et al. [32], we use the vector from $x_{\text{grasp}}$ to $x_{\text{function}}$ to specify the orientation of the object and ask the VLM to predict the orientation from a finite set of options (e.g. forward, backward, upside, downside, left, right).
@@
Since robust grasping relies on contact physics and gripper design, which is beyond the capability of existing pre-trained VLMs, we combine the VLM predictions with analytical approaches in robotic grasping pipelines. Similar to Manuelli et al. [32], we use a grasp sampler to propose candidate grasps based on local geometric information from the observed point cloud. Instead of directly relying on the predicted $x_{\text{grasp}}$, we use the position and orientation of the grasp candidate that is closest to $x_{\text{grasp}}$. For more implementation details, please refer to Appendix B.
@@ IV-E Bootstrapping through Physical Interactions
We consider two ways of further bootstrapping the performance of MOKA through physical interactions with the real world. In both ways, we unroll the VLM policy for the target task to collect robot experiences. The success labels of the collected experiences will be annotated by the VLM or humans.
In-context learning. We use a handful of successful trajectories as in-context examples to guide the high-level and low-level reasoning of the VLM. As discussed in prior work [4], two to three such in-context examples can significantly boost the performance of the VLM. Without changing the prompts $p_{\text{high}}$ and $p_{\text{low}}$, we append the prompt with three pairs of annotated images and VLM responses from previous successful rollouts by the VLM policy.
Policy distillation. Given successful rollouts of the target task, we can train a student policy to imitate the VLM’s successful trajectories. Effectively applying such an approach usually requires data of much larger scales, which often needs substantial efforts in data collection. This approach can showcase how MOKA can be used to collect and bootstrap on real-world robotic datasets for future imitation learning and reinforcement learning algorithms.

### v3

@@ IV-A Point-Based Affordance Representations
To leverage VLMs for solving open-world manipulation tasks, there needs to be an interface that connects the predictions by the VLM and the motions performed by the robot. To achieve this goal, we design an affordance representation defined on 2D images. Produced as the end result by the VLM, the affordance representation specifies the desired motion.
@@
By extending the definitions in Manuelli et al. [42] and Qin et al. [49], we design a point-based affordance representation that can cover a wide range of manipulation tasks. Instead of separately devising motion primitives for different pre-defined skills, we use a unified set of keypoints and waypoints to specify the motion. These points are predicted by VLMs on 2D images and converted to poses in the $SE(3)$ space. Then a smooth motion trajectory is generated based on these poses. To perform the task, the robot interacts with the environment by following the generated motion trajectory.
@@
We specify the robot’s motion in an object-centric manner, as shown in Fig. 2. Following the discussion in Sec. III, we would like this representation to apply to different types of interactions with objects in the environment. Therefore, we consider two types of objects, $o_{\text{in-hand}}$ (e.g., the broom) and $o_{\text{unattached}}$ (e.g., the trash), and specify the motion with a grasping phase and a manipulation phase. In the grasping phase, the robot reaches and grasps an object $o_{\text{in-hand}}$ from the environment. Then in the manipulation phase, the robot performs a motion and makes contact with another object $o_{\text{unattached}}$, either directly or using $o_{\text{in-hand}}$ as a tool. In some scenarios, only one of these two types of objects is interacted with by the robot, either $o_{\text{in-hand}}$ (e.g., unplugging a cable, opening a drawer) or $o_{\text{unattached}}$ (e.g., pressing a button), and one of the two phases can be skipped accordingly.
@@
We now describe the definition of the keypoints and waypoints as well as how they are used to specify the motions in both phases. These points are illustrated in Fig. 2, and more examples can be found in Sec. V-C. Following the practice of Manuelli et al. [42] and Qin et al. [49], we use the grasping keypoint $x_{\text{grasp}}$ to specify the position on $o_{\text{in-hand}}$ where the robot gripper should hold the object. If $o_{\text{in-hand}}$ is not involved in a task, the grasping phase will be skipped. For the manipulation phase, the robot’s gripper follows a motion trajectory specified by an additional set of points. The function keypoint $x_{\text{function}}$ specifies the part of $o_{\text{in-hand}}$ that will make contact with $o_{\text{unattached}}$ in the manipulation phase. If $o_{\text{in-hand}}$ is not specified, $x_{\text{function}}$ will be on the robot gripper, and the contact will directly be made between the robot and $o_{\text{unattached}}$. Correspondingly, the target keypoint $x_{\text{target}}$ is the part of $o_{\text{unattached}}$ that will be contacted by $x_{\text{function}}$ during the manipulation phase. We also introduce the pre-contact waypoints $x_{\text{pre-contact}}$ and the post-contact waypoints $x_{\text{post-contact}}$ defined in free space, which dictates the manipulation motion along with the keypoints defined on the objects.
@@
During the manipulation phase, the robot moves the gripper such that $x_{\text{function}}$ follows the path that sequentially connects the $x_{\text{pre-contact}}$, $x_{\text{target}}$, and $x_{\text{post-contact}}$. Besides following the path, we also require the robot gripper to follow the specified grasping orientation $R_{\text{grasp}}$ and manipulation orientation $R_{\text{manipulate}}$ during the two phases, respectively. To better illustrate the design of our point-based motion, we provide examples of the predicted point specifications and the resultant motions from our experiments in Sec. V-C and the Appendix. In MOKA, this set of keypoints and additional attributes (described in Sec. IV-C) are summarized in a dictionary as the affordance representation (see Fig. 2).
@@ IV-B Affordance Reasoning with Vision-Language Models
To predict the defined affordance representations, we employ the VLM $\mathcal{M}(\cdot)$ (defined in Sec. III), which is pre-trained on Internet-scale data for solving general visual question-answering (VQA) problems. Using a hierarchical prompting framework as shown in Fig. 2, MOKA converts affordance reasoning into a series of VQA problems that are solvable by the pre-trained VLM.
@@
The hierarchical prompting framework takes as input the free-form language description $l$ of the task and an RGB image observation of the environment $s_{t}$. MOKA examines the initial observation $s_{0}$ and decomposes the task $l$ into a sequence of subtasks using the VLM. For each of the subtasks, the VLM is asked to provide the summary of the subtask instruction, the description of the corresponding $o_{\text{in-hand}}$, $o_{\text{unattached}}$, as well as the description of the motion (e.g., “from left to right”). On the low level, given the response from the high-level reasoning and the visual observation $s_{t(k)}$ at the beginning of $k$-th subtask (at the time step $t(k)$), the VLM is queried again with a different prompt to produce the affordance representation defined in Sec. IV-A. In the remainder of this section, we will describe the input formats, output formats, and prompt designs that we use to instantiate this method. Further details can be found in the Appendix.
@@
High-level reasoning. Given the initial observation $s_{0}$ and the language description $l$, we first query the VLM $\mathcal{M}$ with the language prompt $\rho_{\text{high}}$ to produce the response $y_{\text{high}}$:
$y_{\text{high}}=\mathcal{M}([\rho_{\text{high}},l,s_{0}]).$ (1)
@@
The representation $y_{\text{high}}$ is a string that contains structured information for the $K$ subtasks that are needed to solve the task. We design the prompt so as to require the VLM to produce $y_{\text{high}}$ as a list of dictionaries. As shown in Fig. 2, each dictionary contains the language description of a subtask (e.g., “Wipe the snack package to the right side of the table using the broom.”), as well as detailed information to facilitate motion generation, including the description of $o_{\text{in-hand}}$ (e.g., “broom”), the object name of $o_{\text{unattached}}$ (e.g., “snack package”), and the description of the motion (e.g., “from left to right”). This high-level plan will be used as an intermediate result for producing the detailed affordance representation through the low-level reasoning with the VLM. The prompt $\rho_{\text{high}}$ used for high-level reasoning contains the description of the request (the input instruction and the annotated image) and the description of the response format. The complete high-level reasoning prompts are described in TABLE V.
@@
Low-level reasoning. Next, we prompt the VLM once again to produce the affordance representation defined in Sec. IV-A as $y_{\text{low}}^{k}$, conditioning on the high-level representation $y_{\text{high}}$ and the visual observation $s_{t(k)}$ at the beginning of the $k$-th subtask. Instead of directly predicting 3D coordinates on 2D images, which is challenging and even ill-defined, we query the VLM to output 2D coordinates on the images and deproject them back to the 3D space. The three keypoints $x_{\text{grasp}},x_{\text{function}}$ and $x_{\text{target}}$ are defined on the object surface, and thus we can compute the 3D coordinates using the corresponding depth value of the 2D location based on the RGB image and camera parameters. For the waypoints in free space, we query the VLM to predict the desired height in the text. To produce such an affordance representation $y_{\text{low}}^{k}$, we query the VLM again by
$y_{\text{low}}^{k}=\mathcal{M}([\rho_{\text{low}},y_{\text{high}}^{k},f(s_{t(k)})]),$ (2)
where $y_{\text{high}}^{k}$ is the substring corresponding to the $k$-th subtask extracted from $y_{\text{high}}$, and $f(\cdot)$ is a function that process the raw visual observation $s_{t(k)}$. The prompt $\rho_{\text{low}}$ used for low-level reasoning contains the description of the request (the response format of the high-level reasoning and the annotated image) and the description of the response (explanation of the point-based affordance representation and the desired format of the response). The complete low-level reasoning prompts are described in TABLE V, TABLE VI, and TABLE VII.
@@ IV-C Mark-Based Visual Prompting
Inspired by Yang et al. [64], MOKA uses a set of marks as visual prompts to enable VLM to apply its reasoning capability to predict the point-based affordance representation as shown in Fig. 2. Consisting of dots, grids, and text notations annotated on the visual observation, these marks play an important role in the reasoning process. Proposed by open-world object detection and segmentation algorithms, these marks facilitate visual reasoning by encouraging the VLM to attend to the target objects and other task-relevant information in the image. We annotate marks as candidate parts and regions for the VLM to choose the points from, converting the original problem of directly generating coordinates into multiple-choice questions, which is usually more tractable for existing VLMs.
@@
To select the keypoints, which are defined on the in-hand object $o_{\text{in-hand}}$ and the unattached object $o_{\text{unattached}}$ suggested by the high-level reasoning in Sec. IV-B, we propose and plot candidate keypoints on these objects. Given the names of $o_{\text{in-hand}}$ and $o_{\text{unattached}}$, we first segment these two objects using GroundedSAM [54], which combines GroundingDINO [36] and SAM [70] to extract segmentation masks of objects specified by a text prompt. After we obtain the segmentation masks of $o_{\text{in-hand}}$ and $o_{\text{unattached}}$, we perform farthest point sampling [48] on the object contour to obtain $K$ boundary points. Together with its geometric center, and overlay the $K+1$ candidate keypoints on each object. Each candidate keypoint is assigned an index, which is annotated next to it as a reference. To avoid confusion, we use different colors for candidate keypoints on $o_{\text{in-hand}}$ and $o_{\text{unattached}}$ and use the caption in the format of $P_{i}$ and $Q_{j}$ respectively, where $i$ and $j$ are integers. More implementation details can be found in Appendix B-B.
@@
Selecting waypoints in free space involves searching over a much larger region. Instead of directly sampling points in the entire workspace, we divide the observed RGB image into an $m\times n$ grid, where $m$ and $n$ are integers. Both $m$ and $n$ are set to $5$ for our evaluation tasks. The VLM is prompted to choose the tiles in which the pre-contact and post-contact keypoints are supposed to locate in, and then the exact waypoints are sampled uniformly within the tile. For this purpose, we overlay the grid along with the name of each tile on the image. The tile names follow chess notation, which uses letters to specify the columns and integers for the rows.
IV-D Motion Generation with Predicted Affordances
@@
After obtaining the selected keypoints and waypoints from the VLM, we convert the affordance to an executable motion on a real robot. To achieve this goal, we need to lift up all the points from the 2D image to the $SE(3)$ space.
@@
For the keypoints defined on the object, we can directly take the corresponding point on the registered depth image and transform it into the robot’s frame. For the waypoints that are selected from the free space, since they are not attached to any objects, their height needs to be specified in order to be deprojected to 3D space. For this purpose, we also ask the VLM to determine the height of waypoints. We only consider the cases where the waypoints are at the same height as the target point (e.g., pushing) or above the target point (e.g., placing) in most common tabletop manipulation scenarios. Additionally, we also prompt the VLM to return the orientation of $o_{\text{in-hand}}$ during the manipulation phase, which would usually affect the success of tool use tasks.
@@
Drawing inspirations from Manuelli et al. [42], we use the vector from $x_{\text{grasp}}$ to $x_{\text{function}}$ to specify the orientation of the object and ask the VLM to predict the orientation from a finite set of options (e.g., forward, backward, upside, downside, left, right).
@@
Since robust grasping relies on contact physics and gripper design, which is beyond the capability of existing pre-trained VLMs, we combine the VLM predictions with analytical approaches in robotic grasping pipelines. Similar to Manuelli et al. [42], we use a grasp sampler to propose candidate grasps based on local geometric information from the observed point cloud. Instead of directly relying on the predicted $x_{\text{grasp}}$, we use the position and orientation of the grasp candidate that is closest to $x_{\text{grasp}}$. More implementation details can be found in the Appendix.
@@ IV-E Bootstrapping through Physical Interactions
We consider two ways to bootstrap the performance of MOKA through physical interactions with the real world. In both ways, we unroll the VLM policy for the target task to collect robot experiences. The success labels of the collected experiences will be annotated by the VLM or humans.
In-context learning. We use a handful of successful trajectories as in-context examples (as shown in Fig. 7) to guide the high-level and low-level reasoning of the VLM. As discussed in prior work [6], two to three such in-context examples can significantly boost the performance of the VLM. Without changing the prompts $\rho_{\text{high}}$ and $\rho_{\text{low}}$, we append the prompt with three pairs of annotated images and VLM responses from previous successful rollouts by the VLM policy.
Policy distillation. MOKA can be used for collecting demonstration datasets for real-world robot learning. We record the multi-view images and robot proprioception states when collecting successful trajectories and train a visuomotor policy using a simple behavior cloning objective. More details can be found in Sec. V-A.